OpenStudy (anonymous):

cosx=5/13 and sin y=3/5 what is the exact value of cos (x+y)?

OpenStudy (anonymous):

cos(X+y) = cosx*cosy-sinx*siny

OpenStudy (anonymous):

cos(x+y)=cosx.cosy-sinx.siny=5/13*3/5-12/13*4/5

OpenStudy (sriram):

cos x =5/13 hence sin x=12/13 (using sin^2x + cos^2 x=1) sin y=3/5 cos y=4/5 cos (x+y)=cos x*cos y - sin x*sin y =5/13 *4/5 - 12/14 *3/5 =20/65-36/65 =(20-36)/65 =-16/65